OpenStudy (anonymous):

if the solution of (x+1)(3x^2-2x-1)=0 then x+1/x= ?

OpenStudy (anonymous):

x = -1, -1/3, 1

OpenStudy (anonymous):

(x+1) (3x+1) (x-1) = 0 x+1=0 x=-1 sub to x+1/ x =0/-1 = 0 3x+1=0 x=-1/3 sub to x+1/ x = (2/3) / (-1/3) = -2 x-1 =0 x=1 sub to x+1/ x = 2/1 = 2